OpenStudy (anonymous):

A factory can produce two products,x and y, with a profit approximated by P=14x+22y-900. The production of y must exceed the production of x by at least 100 units. Moreover, production levels are limited by the formula x+2y≤1400. a. Identify the vertices of the feasible region. b. What production levels yield the maximum profit, and what is the maximum profit?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

from the text of your problem, we have two inequalities, namely: \[x+2y \le 1400, y \ge 100+x\]

OpenStudy (anonymous):

ok, I didn't get this when it was covered over class...

OpenStudy (michele_laino):

please, note that if: "The production of y must exceed the production of x by at least 100 units" I understand that y>=100+x, do you agree?

OpenStudy (anonymous):

yes that makes sense

OpenStudy (michele_laino):

furthermore, both x and y are positive quantities

OpenStudy (anonymous):

yeah, so how do i find the vertices and max production?

OpenStudy (michele_laino):

so your region is defined by the subsequent inequalities: y<=700-x/2, y>=x+100 x>=0 y>=0

OpenStudy (michele_laino):